# AgentAda: Skill-Adaptive Data Analytics for Tailored Insight Discovery

### v1

@@ 6.3 Evaluating the performance of Skill Matcher
6.4 Ablation: Influence of Goal and Persona
| Depth of Analysis | 19 | 8 | 73 | 0 |
| Relevance To Goal | 75 | 6 | 18 | 1 |
| Persona Consistency | 31 | 13 | 54 | 2 |
| Coherence | 18 | 13 | 67 | 2 |
| Plot Conclusion | 11 | 2 | 86 | 1 |
@@ 7 Conclusion & Future Works
We recruited human evaluators for a 1-hour annotation task and compensated them with a $10 gift card. All evaluators provided informed consent prior to participation and were informed about the nature of the task. The evaluation process adhered to ethical research guidelines, ensuring voluntary participation and fair compensation. No personally identifiable information was collected.
@@ Appendix C Human Evaluation Platform
Depth of Analysis: Looks at how deeply the data was explored and whether meaningful insights were uncovered.
@@
Relevance to Goal: Checks if the analysis stays focused on what the project set out to achieve.
@@
Persona Consistency: Sees if the work is tailored to the intended user’s needs, background, and expectations.
@@
Coherence: Evaluates how smoothly and logically the ideas and findings are connected throughout the analysis.
@@
Model Adequacy: Identifies the model that best solves the main problem or meets the objective.
@@
Which one shows a proper conclusion of the plot: Determines which plot or result clearly wraps up the analysis with a solid takeaway.
@@ Appendix D Human Evaluation Statistics and Details
We recruited 30 participants through a Google Form, which included task instructions and an estimated completion time based on our pilot study (1.5–2.5 minutes). Among the participants, 21 were male and 9 were female. As detailed in C, we also recorded each participant’s professional designation. Figure 9 shows the distribution of evaluator expertise.
Appendix E Prompts
E.1 KaggleBench Prompts
@@
Prompt 1 and Prompt 2 illustrates the prompts used for generating the question answer pairs and goal & persona for KaggleBench respectively.
E.2 Dual Stage Advanced Question Generation Prompts
@@
To generate dataset-specific questions, we initially prompted GPT-4o-mini (OpenAI et al., 2024) to produce five basic questions that aid data analysts in understanding a dataset. The prompt (see Prompt 3) accepts input parameters such as the dataset’s analysis goal, the analyst’s persona, the names and data types of the dataframe columns, and the dataframe head. An output template is also provided to ensure consistent formatting. The primary objective of this prompt is to generate five questions that offer fundamental insights into the dataset.
@@
Subsequently, these basic questions—along with the original input information—are fed into a specialized advanced question generation prompt (see Prompt 4). This prompt, also leveraging GPT-4o-mini (OpenAI et al., 2024), is designed to generate skill-oriented questions. We supply an output format template that organizes the output into distinct task and question components for consistency. The main focus of this advanced prompt is to produce questions that require the advanced analytical skills defined in our skill library, thereby uncovering deeper insights into the dataset and yielding more actionable results.
@@
We also explored a single-prompt approach for the question generation pipeline. The prompt (see Prompt 5) accepts the same inputs as our advanced question generation prompt, with the exception of the basic generated questions. However, this approach yielded questions that were either overly similar or did not align with the advanced analytical requirements we aimed to address.
@@
E.3 Category Prediction Prompts
To guide GPT-4o in predicting high-level insight themes for a given dataset, we design a structured prompt that provides the model with (I) the dataset description, (II) the overall analysis goal, and (III) the list of generated analytical questions. The goal of the prompt is to predict exactly three distinct, meaningful categories that are broad enough to group multiple related insights but specific enough to remain actionable and aligned with the context of the analysis. The prompt (see Prompt 6) emphasizes:
@@ E.3 Category Prediction Prompts
E.4 Skill matcher Prompts
To identify the most relevant skills for a given question, we prompt GPT-4o (OpenAI et al., 2024) with the question and a list of all available skills in the library. The prompt asks the model to rank the top three skills based on their usefulness in answering the question. We also provide a structured output template to ensure consistency in formatting. Refer to Prompt 7 for more details.
E.5 Code Generation Prompts
To generate the required plot for answering a question, we prompt GPT-4o (OpenAI et al., 2024) using the question and a summary of the selected skill. The prompt (see Prompt 8) is responsible for generating both the code and key statistics about the dataset. It emphasizes structured code generation, producing code that encompasses data preparation, skill application, visualization, computation of key statistics, and adherence to best coding practices.
To ensure that the generated code utilizes the required skill, we pass the code along with our skill list to GPT-4o for verification. This check is performed using Prompt 9.
E.6 Answer Generation Prompts
For each question, we execute the code generated in Appendix E.5 to obtain statistics and plots. These outputs serve as multimodal inputs to GPT-4o, which extracts answers using Prompt 10. This prompt is designed to identify key patterns, anomalies, comparisons, and notable findings from both the visualizations and statistics, capturing all relevant qualitative and quantitative details. Subsequently, the answer summarizer prompt (see Prompt 11) condenses these findings to the top two key points, producing concise, single-line answers that are supported by quantitative evidence.
E.7 Insight Generation Prompts
The individual answers are aggregated to derive key observations and actionable insights for the entire dataset. Prompt 12 leverages the curated answers, along with the predicted categories from Appendix E.3, the analysis goal, and the dataset description, to generate the final insights. This prompt focuses on distilling the most critical and meaningful insights, ensuring that they are presented in a structured format and backed up by quantitative evidence.
Appendix F SCORER
@@
Appendix G Qualitative Analysis
@@
Appendix H Task-wise Human Evaluation Results
@@
Appendix I Detailed Question-wise LLM Evaluation Results
Appendix J Detailed Insight-wise LLM Evaluation Results

### v2

@@ 7 Conclusion & Future Works
Our work on AgentAda does not involve personal or sensitive data, but it raises important considerations. All datasets used are curated from public or sanitized sources, and all human annotators gave informed consent and were fairly compensated. In real-world deployments, however, analytics agents may interact with proprietary or sensitive data, so safeguards such as access control, anonymization, and auditing are essential. Moreover, automatic method selection can inadvertently propagate bias in the data, underscoring the need for fairness audits and transparency in reasoning.
We employed Large Language Models (LLMs) only as writing aids to improve clarity, grammar, and readability of the manuscript. The conceptual framing, methodological design, experimental implementation, analysis, and conclusions remain entirely the work of the authors.
Reproducibility Statement
We have taken several steps to ensure reproducibility. The AgentAda codebase, skill library, and evaluation scripts will be released under a permissive license. KaggleBench benchmark will also be made publicly available. Implementation details, prompt templates, hyperparameters, and ablation settings are documented in the appendices, and each analytical skill in the library is accompanied by clear input/output documentation and usage examples. Finally, we release baseline agents and evaluation protocols to replicate our experiments and results, making the system transparent and extensible for future research.
@@ Appendix C Human Evaluation Platform
Depth of Analysis: Evaluates whether the response goes beyond surface-level descriptions to uncover non-trivial patterns, relationships, or trends in the data. A high score indicates multi-step reasoning and interpretation that adds real analytical value.
@@
Relevance to Goal: Assesses how well the response remains aligned with the stated analytical objective or research question. Strong responses avoid unnecessary detours and keep the analysis tightly focused on achieving the intended goal.
@@
Persona Consistency: Measures whether the response is consistent with the intended persona’s expertise, perspective, and communication style. For instance, a “business analyst” persona should emphasize actionable insights, while a “data scientist” persona should highlight methodological rigor.
@@
Coherence: Examines the clarity, logical structure, and internal consistency of the response. Higher scores indicate that arguments, evidence, and conclusions are presented in a well-organized and easy-to-follow manner.
@@
Answering Question Adequately: Determines whether the response fully and correctly addresses the question posed. Partial answers, misinterpretations, or omissions lower the score, while comprehensive and precise responses score higher.
@@
Plot Conclusion Quality: Evaluates the correctness and clarity of conclusions drawn from plots or visualizations. Strong responses explicitly connect the visual evidence to the broader analysis, offering a clear and accurate takeaway rather than vague or generic statements.
@@ Appendix D Human Evaluation Statistics and Details
We recruited 30 participants through a Google Form, which included task instructions and an estimated completion time based on our pilot study (1.5–2.5 minutes). Among the participants, 21 were male and 9 were female. As detailed in Appendix C, we also recorded each participant’s professional designation. Figure 9 shows the distribution of evaluator expertise.
| 0.8842 | 0.8297 | 0.8431 | 0.7658 | 0.8274 | 0.8765 |
Appendix E Ablation: Comparison with other code generators
We next compare AgentAda with two representative code generation baselines ECA: (Wang et al., 2024) and Data Interpreter (Hong et al., 2024). Both approaches rely on invoking existing tools or examples but lack the structured, skill-based generation that distinguishes AgentAda.
ECA: The prompt used for ECA is given in Prompt 1. ECA generates code by directly calling statistical or machine learning tools (e.g., ARIMA, Granger causality tests) without constructing full pipelines. As a result, it frequently misses prerequisite steps such as preprocessing, diagnostics, or lag order selection, leading to incomplete or misleading analyses. For example, in the Granger causality task in the Table 9, ECA reduces the problem to testing a variable against its own lag, producing an ninformative setup. This reflects a fundamental limitation, ECA is a tool invoker rather than an analysis engine.
Data Interpreter (MetaGPT). Data Interpreter generates code using few-shot prompting, retrieving semantically similar examples and directly stitching together function calls. While this can yield runnable code, it often fails to produce task-specific pipelines aligned with the analytical goal, since it does not incorporate user intent or persona. By contrast, AgentAda’s access to skill library act as structured blueprints that allow AgentAda to generate code from scratch, personalized to the goal and persona.
AgentAda. The key distinction is that AgentAda is not a code generator but a full analysis engine. By grounding generation in skills, it produces pipelines entirely from scratch using skills from the skill library that encapsulate step-by-step workflow. Crucially, this is not borrowed or stitched-together code, but code that is personalized to the specific task which ensures necessary preprocessing, diagnostics, modeling, and evaluation steps are correctly integrated.
We also conducted a quantitative comparison of AgentAda, ECA, and EvoR on all QA pairs of KaggleBench. Accuracy was measured directly against the benchmark QA answers, while rubric-based dimensions were judged on a 1–5 scale using GPT-4o tuned with SCORER. The rubrics evaluate depth of analysis, adequacy of answers, coherence, relevance to goal, and persona consistency.
As shown in Table 10, AgentAda achieves the highest accuracy (90.6%) and outperforms both ECA and EvoR (Su et al., 2024b) across all rubric dimensions. These results are consistent with the above qualitative case study on Granger causality where AgentAda generated a complete and methodologically sound pipeline. Together, the quantitative and qualitative evidence highlight AgentAda’s advantage as a full analysis engine that generates task-specific, coherent, and goal-aligned analyses, unlike baselines that rely on tool invocation or code stitching.
| Model | Accuracy | Depth of Analysis | Answers Adequately | Coherence | Relevance to Goal | Persona Consistency |
| AgentAda | 90.6 | 4.46 | 4.41 | 4.48 | 4.44 | 4.42 |
| ECA | 86.4 | 4.15 | 3.96 | 4.34 | 4.12 | 3.86 |
| EvoR | 86.8 | 4.22 | 4.03 | 4.39 | 4.18 | 3.92 |
Appendix F Evaluating the Factuality of the Insights
We further assess the factual consistency of agent outputs using multiple factuality metrics. For large proprietary models such as GPT-5 OpenAI (2025), GPT-4o OpenAI et al. (2024) LLaMA-3.3-70B Grattafiori et al. (2024), we directly query the models to provide a factuality rating on a – scale, where 1 indicates that the output contains many factual errors and is mostly ungrounded, while 5 indicates that the output is fully consistent with the provided data and free of factual mistakes. In addition, we employ FactScore Min et al. (2023), which evaluates factuality by decomposing model outputs into atomic claims and verifying them against ground-truth sources. In our setup, we follow the original FactScore methodology and use the KaggleBench dataset, consisting of 700 examples across diverse notebooks. For each notebook, we utilize the associated QA list to extract evaluation questions and restrict atomic claims to those grounded in the underlying data tables, excluding suggestions or speculative outputs not present in the source. We compute FactScore with multiple verifiers, including GPT-5, GPT-4o, LLaMA-3.3-70B, and Claude-Opus 4.
Table 11 reports the results. We observe that AgentAda with skills consistently outperforms its ablation and other baselines across all factuality metrics, confirming that skills contribute to generating more accurate, data-grounded insights.
| Model | LLM as a Judge | FactScore | ||||
| GPT-5 | GPT-4o | LLaMA-3.3-70B | GPT-5 | GPT-4o | LLaMA-3.3-70B | |
| AgentAda w skill | 4.30 | 4.25 | 4.12 | 87.9 | 88.1 | 87.4 |
| AgentAda w/o skill | 3.85 | 3.82 | 3.71 | 76.1 | 76.3 | 75.8 |
| MetaGPT | 3.98 | 3.95 | 3.84 | 82.1 | 82.4 | 81.9 |
| InfiAgent | 3.92 | 3.90 | 3.79 | 80.0 | 80.1 | 79.7 |
| Poirot | 3.80 | 3.78 | 3.65 | 73.7 | 73.9 | 73.3 |
| Pandas | 3.76 | 3.75 | 3.62 | 71.6 | 71.8 | 71.2 |
| GPT-4o | 3.74 | 3.70 | 3.59 | 72.7 | 72.9 | 72.5 |
Appendix G Ablation: Influence of Goal and Persona
| Depth of Analysis | 19 | 8 | 73 | 0 |
| Relevance To Goal | 75 | 6 | 18 | 1 |
| Persona Consistency | 31 | 13 | 54 | 2 |
| Coherence | 18 | 13 | 67 | 2 |
| Plot Conclusion | 11 | 2 | 86 | 1 |
Appendix H Ablation: Alignment of SCORER with Human Feedback
In this section we evaluate how well different evaluators align with human judgment when comparing two variants of our system: AgentAda with skills versus AgentAda without skills. Each evaluator must decide which variant produces better insights on KaggleBench datasets, or mark the outcome as a tie or none. We consider four evaluators: SCORER, Human Evaluation, G-Eval, and LLaMA-Eval. Note that, the data used to train SCORER is disjoint from the validation set used for this experiment, ensuring no leakage or bias. As shown in Table 13, SCORER aligns much more closely with human evaluation than G-Eval and LLaMA-Eval. This is expected, as SCORER is explicitly optimized for human alignment.
| Evaluation | ADA w/ Skill Win | Tie | w/o Skill Win | None |
| SCORER | 53.3 | 22.6 | 22.3 | 1.8 |
| Human Evaluation | 44.5 | 27.2 | 25.9 | 2.4 |
| G-Eval | 28.7 | 41.2 | 27.5 | 2.6 |
| LLaMA-Eval | 28.2 | 43.0 | 26.1 | 2.6 |
Appendix I Prompts
I.1 KaggleBench Prompts
@@
Prompt 2 and Prompt 3 illustrates the prompts used for generating the question answer pairs and goal & persona for KaggleBench respectively.
I.2 Dual Stage Advanced Question Generation Prompts
@@
To generate dataset-specific questions, we initially prompted GPT-4o-mini (OpenAI et al., 2024) to produce five basic questions that aid data analysts in understanding a dataset. The prompt (see Prompt 4) accepts input parameters such as the dataset’s analysis goal, the analyst’s persona, the names and data types of the dataframe columns, and the dataframe head. An output template is also provided to ensure consistent formatting. The primary objective of this prompt is to generate five questions that offer fundamental insights into the dataset.
@@
Subsequently, these basic questions—along with the original input information—are fed into a specialized advanced question generation prompt (see Prompt 5). This prompt, also leveraging GPT-4o-mini (OpenAI et al., 2024), is designed to generate skill-oriented questions. We supply an output format template that organizes the output into distinct task and question components for consistency. The main focus of this advanced prompt is to produce questions that require the advanced analytical skills defined in our skill library, thereby uncovering deeper insights into the dataset and yielding more actionable results.
@@
We also explored a single-prompt approach for the question generation pipeline. The prompt (see Prompt 6) accepts the same inputs as our advanced question generation prompt, with the exception of the basic generated questions. However, this approach yielded questions that were either overly similar or did not align with the advanced analytical requirements we aimed to address.
@@
I.3 Category Prediction Prompts
To guide GPT-4o in predicting high-level insight themes for a given dataset, we design a structured prompt that provides the model with (I) the dataset description, (II) the overall analysis goal, and (III) the list of generated analytical questions. The goal of the prompt is to predict exactly three distinct, meaningful categories that are broad enough to group multiple related insights but specific enough to remain actionable and aligned with the context of the analysis. The prompt (see Prompt 7) emphasizes:
@@ I.3 Category Prediction Prompts
I.4 Skill matcher Prompts
To identify the most relevant skills for a given question, we prompt GPT-4o (OpenAI et al., 2024) with the question and a list of all available skills in the library. The prompt asks the model to rank the top three skills based on their usefulness in answering the question. We also provide a structured output template to ensure consistency in formatting. Refer to Prompt 8 for more details.
I.5 Code Generation Prompts
To generate the required plot for answering a question, we prompt GPT-4o (OpenAI et al., 2024) using the question and a summary of the selected skill. The prompt (see Prompt 9) is responsible for generating both the code and key statistics about the dataset. It emphasizes structured code generation, producing code that encompasses data preparation, skill application, visualization, computation of key statistics, and adherence to best coding practices.
To ensure that the generated code utilizes the required skill, we pass the code along with our skill list to GPT-4o for verification. This check is performed using Prompt 10.
I.6 Answer Generation Prompts
For each question, we execute the code generated in Appendix I.5 to obtain statistics and plots. These outputs serve as multimodal inputs to GPT-4o, which extracts answers using Prompt 11. This prompt is designed to identify key patterns, anomalies, comparisons, and notable findings from both the visualizations and statistics, capturing all relevant qualitative and quantitative details. Subsequently, the answer summarizer prompt (see Prompt 12) condenses these findings to the top two key points, producing concise, single-line answers that are supported by quantitative evidence.
I.7 Insight Generation Prompts
The individual answers are aggregated to derive key observations and actionable insights for the entire dataset. Prompt 13 leverages the curated answers, along with the predicted categories from Appendix I.3, the analysis goal, and the dataset description, to generate the final insights. This prompt focuses on distilling the most critical and meaningful insights, ensuring that they are presented in a structured format and backed up by quantitative evidence.
Appendix J SCORER
@@ I.7 Insight Generation Prompts
Appendix K AgentAda Generalizability
To evaluate the generalizability of our approach, we measured the performance of AgentAda and compared it with other baseline agents across two distinct benchmarks: InsightBench (Sahu et al., 2024) and InfiAgent-DABench (Hu et al., 2024). As shown in Table 14, AgentAda with GPT-4o consistently achieves the highest performance, attaining the best scores in both evaluation settings. Importantly, when replacing GPT-4o with alternative backbones such as LLaMA-3.3-70B (Grattafiori et al., 2024) or Qwen-Coder-2.5-7B (Hui et al., 2024; Team, 2025), AgentAda remains competitive and continues to outperform code generation methods that do not leverage a skill library. This demonstrates that the pipeline is robust not only across datasets but also across different underlying models. Furthermore, other baselines such as InfiAgent, MetaGPT, and Poirot show a clear gap compared to AgentAda, with performance dropping significantly across both benchmarks. These results highlight that our design generalizes effectively to diverse tasks and architectures, while maintaining a substantial margin over existing agents.
| Agent | InsightBench | InfiAgent-DABench | ||
| G-Eval | LLaMA3-Eval | Accuracy | F1 | |
| AgentAda (GPT-4o) | 81.7 | 83.2 | 76.5 | 78.3 |
| AgentAda (LLaMA-3.3-70B) | 79.8 | 80.5 | 73.6 | 75.1 |
| AgentAda (Qwen-Coder-2.5-7B) | 78.9 | 79.2 | 72.8 | 74.0 |
| AgentAda (Code-Llama-Instruct-13B) | 75.4 | 76.1 | 69.2 | 70.8 |
| w/o Skill (GPT-4o) | 72.3 | 71.7 | 66.2 | 67.8 |
| InfiAgent | 69.8 | 68.4 | 62.0 | 63.7 |
| MetaGPT | 67.5 | 66.1 | 60.1 | 61.9 |
| Poirot | 63.4 | 62.2 | 58.9 | 59.3 |
| GPT-4o | 61.7 | 60.3 | 55.2 | 56.8 |
| Pandas | 59.1 | 57.6 | 52.7 | 53.5 |
Appendix L SCORER Generalizability
Beyond demonstrating the effectiveness of introduced evaluation mechanism, SCORER, we also evaluate the robustness of it. We investigate two key aspects of generalizability: (1) robustness to different evaluation models, including open-source alternatives, and (2) applicability to new datasets where SCORER was not explicitly optimized.
Table 15 shows that SCORER generalizes well across evaluation models. For example, on InsightBench, SCORER using GPT-4o judges yields 43.2% wins for AgentAda with skills versus 15.6% for the version without skills, while LLaMA-3.3-70B judges give a nearly identical result (42.3% vs. 15.3%). Similar patterns hold for DS-1000 and KaggleBench, demonstrating that SCORER remains consistent regardless of the underlying evaluation model. This confirms its utility for supporting open-source evaluation pipelines.
| Dataset | SCORER Judge Model | ADA w Skill Win | Tie | ADA w/o Skill Win | None |
| InsightBench | GPT-4o | 43.2 | 40.7 | 15.6 | 0.5 |
| LLaMA-3.3-70B | 42.3 | 41.8 | 15.3 | 0.6 | |
| DS-1000 | GPT-4o | 42.8 | 41.0 | 15.5 | 0.7 |
| LLaMA-3.3-70B | 42.9 | 41.2 | 15.4 | 0.5 | |
| KaggleBench | GPT-4o | 53.3 | 22.6 | 22.3 | 1.8 |
| LLaMA-3.3-70B | 52.5 | 25.4 | 20.8 | 1.3 |
We also evaluate SCORER on benchmarks where it was not specifically tuned. As shown in Table 16, SCORER provides results that align with standard absolute metrics such as G-Eval, BERTScore, and ROUGE-L. For example, on InsightBench, AgentAda with skills achieves 81.7 G-Eval compared to 72.3 without skills, and SCORER correspondingly shows more wins for the skill-based version (33.2% vs. 12.6%) alongside a high number of ties. A similar trend is observed on DS-1000, further validating SCORER’s consistency. Unlike purely semantic metrics such as BERTScore, SCORER aligns more closely with human judgment by capturing fine-grained improvements due to skill integration, demonstrating its generalizability across datasets.
| Dataset | Relative Scoring | Absolute Scoring | ||||||||
| SCORER | AgentAda w Skill | AgentAda w/o Skill | ||||||||
| w Skill Win | Tie | w/o Skill Win | None | G-Eval | BERTScore | ROUGE-L | G-Eval | BERTScore | ROUGE-L | |
| InsightBench | 33.2 | 53.7 | 12.6 | 0.5 | 81.7 | 85.2 | 44.7 | 72.3 | 81.4 | 40.3 |
| DS-1000 | 32.8 | 55.1 | 11.4 | 0.7 | 74.0 | 82.7 | 42.1 | 67.8 | 78.2 | 38.7 |
Appendix M Qualitative Analysis
@@
Appendix N Task-wise Human Evaluation Results
@@
Appendix O Detailed Question-wise LLM Evaluation Results
Appendix P Detailed Insight-wise LLM Evaluation Results